Build a voltaic cell with one beaker containing potassium permanganate (KMnO4 ) solution with a platinum electrode and a second beaker containing zinc chloride with a zinc electrode. Determine:

a. The half reactions

b. Flow of electrons

c. Short-form notation

d. Electric potential (voltage)


1
Expert's answer
2021-01-12T07:00:07-0500

Solution.

a.

(A): "Zn^{2+} + 2e = Zn^0"

(K): "MnO4^- + 2 H2O + 3e = MnO2 + 4OH^-"

b.

Electrons move from the anode (zinc) to the cathode (platinum).

c.

"Zn(s)|Zn^{2+}(aq)||MnO4^-(aq)|Pt(s)"

d.

"\Delta E = E(k)-E(a)"

"\Delta E = 0.6 - (-0.763) = 1.363 \ V"

Answer:

a.

(A): "Zn^{2+} + 2e = Zn^0"

(K):"MnO4^- + 2 H2O + 3e = MnO2 + 4OH^-"

b.

Electrons move from the anode (zinc) to the cathode (platinum).

c.

Zn(s)|Zn^{2+}(aq)||MnO4^-(aq)|Pt(s)

d.

"\Delta E = 1.363 \ V"